OpenStudy (anonymous):

Calculate sum : 1/(n+1) + 1/(n+2)........to 3n terms

OpenStudy (anonymous):

so the rth term is 1/(n+r)

OpenStudy (nikvist):

\[\sum\limits_{i=1}^{2n}\frac{1}{n+i}\] is that correct sum

OpenStudy (anonymous):

oh lodu...wts the answer?

OpenStudy (nikvist):

last term in sum is 1/3n or 1/4n?

OpenStudy (anonymous):

according to me the answer is \[\ln 4\]

OpenStudy (anonymous):

last term is 1/4n

OpenStudy (watchmath):

Do we take the limit as \(n\to \infty\). Otherwise I don't know how to do that.

OpenStudy (anonymous):

thanks metarock

OpenStudy (anonymous):

yes i have taken n tending to infinity and made a definite integral ..